1) An investor was originally expecting a 16% return on her portfolio with beta of 1.25 before
the market risk premium decreased from 8% to 6%. Given this change, what return will
now be expected on the portfolio?

2) John PLC is financed by debt and equity. The market value of its debt is £8 million and
its equity £12 million. If John’s cost of debt is 6% and the required rate of return on
equity is 12%, its Weighted Average Cost of Capital (WACC) is (assume there is no tax):
1
Expert's answer
2018-12-24T09:56:11-0500

1)

We can use CAPM formula:

r_e=r_f+β(r_m-r_f)

where

rf –risk-free rate

rm- market return

β- stock Beta

(r_m-r_f) - market risk premium


From this equation we will find risk-free rate:

r_f=r_e-β*(r_m-r_f )=0.16-1.25*0.08=0.06

Expected return after decreasing risk premium is following:

r_e=0.06+1.25*0.06=0.135= 13.5%


2)

WACC=E/(D+E) (r_e )+D/(D+E) (r_d )

where

E- market value of equity

D – market value of debt

re – cost of equity

rd – cost of debt

WACC=12/(8+12)*0.12+8/(8+12)*0.06=0.096=9.6%